# The Influence of Exclusion Zones on the Coexistence of Predator and Prey with an Allee Effect

### v1

@@
Abstract.
We propose a reaction–diffusion model of predator–prey interaction in which the predators occupy only a subset of the prey’s territory, leaving a predator-free exclusion zone. Ecological examples include marine protected areas where it is illegal to fish, or buffer zones left between the territories of rival predators. The prey are subject to a strong Allee effect, so excessive predation may lead to the extinction of both species. The exclusion zone mitigates this problem by providing the prey with a refuge in which to proliferate without predation. Thus, paradoxically, a smaller predator territory may be able to support a more substantial population than a larger one. Using a topological degree argument, we show in any dimensions that, provided the exclusion zone is large enough, the system possesses spatially heterogeneous coexistence equilibria with positive populations of both species. This result is global in the sense that it does not rely on local bifurcations from semi-trivial stationary states. We also show that as the predator domain becomes asymptotically small, the total predator population does not vanish, and in some cases may actually be maximized in this limit of shrinking predation area. Conversely, we show that as the predator domain becomes large, it may exhibit thresholding behavior, passing suddenly from a regime with coexistence solutions to one in which extinction becomes unavoidable, highlighting the need for careful analysis in the management of predator–prey systems.
Key words and phrases:
reaction–diffusion system, predator–prey model, Allee effect, topological degree, exclusion zone
2020 Mathematics Subject Classification:
Primary 35K57; Secondary 35J57, 35B40, 92D25
1. Introduction
1.1. The role of exclusion zones in predator-prey systems
@@
A somewhat related phenomenon is to be found in the spatial organization of competing (i.e., mutually hostile) groups of predators. For example, groups of predators such as packs of wolves or coyotes often divide available space into sharply defined and defended regions, leading to territory formation [38]. As a consequence of this division, the interfaces between territories constitute buffer zones into which predators avoid adventuring and where prey thrive [29]. It was shown [6] that such territoriality can be beneficial for the overall predator population in spite of casualties from the intraspecific hostility. Thus, in practice, the buffer zones between territories act as exclusion zones in which predators are excluded and prey can proliferate without predation.
@@
1.2. The Predator-Prey model with exclusion zones
@@
$\begin{cases}u_{t}-d_{u}\Delta u=f(u)-\beta\mathbb{1}_{A}uv&x\in\Omega\\ v_{t}-d_{v}\Delta v=-\gamma v+\alpha uv&x\in A\\ u_{x}(x,t)=0&x\in\partial\Omega\\ v_{x}(0,t)=0&x\in\partial A.\end{cases}$ (1)
where $\beta$ is the predation rate, $\gamma$ is the mortality rate of predators in the absence of prey, $\alpha$ is a conversion efficiency, $d_{u}$ and $d_{v}$ are diffusion coefficients of the prey and predator, respectively, $L$ is the domain size, and $a$ is the location of the boundary of the exclusion zone within $L$. Moreover, $A\subset\Omega\subset\mathbb{R}^{N}$ are open, $\alpha,\beta,\gamma>0$, and $f:[0,1]\to\mathbb{R}$ is a smooth function with the following properties. There are exactly three roots of $f$: 0, 11, and $\theta\in(0,1)$ with $f(x)<0$ for $x\in(0,\theta)$ and $f(x)>0$ for $x\in(\theta,1)$. We consider only solutions $u$ and $v$ so that $0\leq u(t,x)\leq 1$ for all $x\in\Omega$ and $0\leq v(t,x)$ for all $x\in A$. Note that these definitions stipulate that $\Omega\setminus A$ is the predator-free exclusion zone (equivalently, the MPA) and that both predators and prey may occur in $A$; this will later prove notationally convenient. Figure 1 shows an example of $\Omega$ and $A$.
Figure 1. A diagram of $A\subset\Omega$ illustrating the assumptions of Theorem 1.4.
As a first result about the model (1), we prove the following:
Proposition 1.1.
If $\alpha<\gamma$, then all solutions to (1) with $0\leq u\leq 1$ and $v\geq 0$ satisfy $\lim_{t\to\infty}\|v\|_{L^{1}}$.
@@
Since the analysis shows that predators inevitably go extinct when $\alpha<\gamma$, this parameter regime is not biologically relevant for sustained predator-prey interactions, and we therefore restrict attention to the case $\gamma\leq\alpha$.
Sections 1.3-5 focus on analytical results for solutions of this model. Much of this analysis concerns equilibrium solutions to (1). Clearly, there are three homogeneous equilibria, namely $v=0$ while $u\in\{0,\theta,1\}$. But we are most interested in inhomogeneous coexistence equilibrium solutions to (1) in which $0<u<1$ and $v>0$. In Sections 1.3 and 1.4, we present the main existence results of this paper, showing that such coexistence equilibria exist in both one and higher dimensions. We prove the existence of coexistence equilibria in one dimension in Section 2 and generalize this argument to higher dimensions in Section 3. The resulting existence theorems include as a condition that the predator-free domain $\Omega\setminus A$ must be sufficiently large. In Section 4, we demonstrate how coexistence equilibria may cease to exist when this condition does not hold. We perform asymptotic analyses of inhomogeneous equilibria as $A$ becomes small (so-called “thin-limit” analysis) in one dimension in Section 5.
@@
  • What is the size (or shape) of the predator domain $A$ that maximizes the predator population?
@@
1.3. Coexistence solutions in the 1D problem
One of the first questions we may ask about the model (1) is about the existence of coexistence equilibria in the 1D case where $\Omega=(0,L)$ and $A=(0,a)$ for $a<L$. That is, we are interested in solutions $u(x),v(x)$ of
[…]
with $v\not\equiv 0$.
To state the result, we consider the auxiliary scalar problem where we look for a function $\zeta$ defined on $[a,L]$:
$\begin{cases}-d_{u}\zeta_{xx}=f(\zeta)&x\in(a,L)\\ \zeta(a)=0,&\zeta_{x}(L)=0.\end{cases}$ (3)
We consider $\zeta$ to be the “worst-case scenario” solution for the prey, modeling a situation where the prey items are killed immediately upon diffusing from the predator-free domain $(a,L)$ into the predator domain at $x=a$. We therefore expect any biologically meaningful coexistence equilibrium $(u,v)$ to (2) to satisfy $u\geq\zeta$, and indeed in Theorem 1.3, we show that a nontrivial solution $\zeta$ of (3) is a subsolution for $u$.
Proposition 1.2.
For b:=Lab:=L-a sufficiently large, equation (3) admits a maximum positive solution $\zeta$ such that $0<\zeta<1$.
This result is classical [47, 24] and follows from standard arguments. Namely, first, $\overline{\zeta}(x)\equiv 1$ is a supersolution. Then, for $b\geq b_{1}$, for some $b_{1}>0$, there exists a solution $0<\underline{\zeta}<1$ of $\zeta_{xx}+f(\zeta)=0$ in $(a,L)$, with $\underline{\zeta}(a)=\underline{\zeta}(L)=0$. This function $\underline{\zeta}<\overline{\zeta}=1$ serves as a sub-solution for problem (3). Thus, by sub- and super-solution principle, there exists a maximum positive solution $\zeta$. Note that $b_{1}$ only depends on LaL-a, not on $a$. In fact, it is also known that there is a threshold value $b^{\ast}$ for b=Lab=L-a such that no positive solution exists when $b\leq b^{\ast}$. This is related to the critical size of domains for equations with Allee effect and absorption on (part of) the boundary.
We can now state our result for non-trivial coexistence solutions in the one-dimensional setting.
Theorem 1.3.
Consider the system (2) with $f$ of bistable type satisfying the previous assumptions. Assume that $\gamma<\alpha$. There exists $b_{0}$ such that for any $a,L$ such that $L-a\geq b_{0}$, there exists a solution $(u,v)$ of system (2) such that $\zeta(x)<u(x)<1$ for all $x\in[a,L]$ and $v>0$ in $[0,a]$. In particular, $u$ is non constant.
This theorem, which we prove in Section 2, gives the existence of coexistence equilibria provided that (1) the predator death rate $\gamma$ is less than the birth rate per prey item $\alpha$, and (2) the predator-free zone $[a,L]$ is sufficiently large. Thus, a coexistence equilibrium exists if the predators can reproduce fast enough and the prey have enough space to proliferate without being hunted. Moreover, the prey density of this coexistence equilibrium exceeds the “worst-case scenario” prey density $\zeta$.
The condition $L-a\geq b_{0}$ cannot be dropped in general, as we shall see in Section 4 if $a=L$ (more generally, if $A=\Omega$) and $\gamma/\alpha<\theta$, then (1) has no coexistence equilibria. In this case, the predators are too successful for their own good, and there is nowhere that the prey are safe. The predators drive the prey beneath the Allee threshold and to extinction, eventually leading to their own extinction as well.
1.4. Coexistence solutions in higher dimensions
Let $\Omega$ be a bounded and smooth domain of $\mathbb{R}^{N}$, the cases of interest in modeling being $N=2,3$. We consider the predator-prey model, where predation is limited to a predation zone $A\subset\Omega$. We assume $A$ to be a smooth subset (connected or not) of $\Omega$. We assume here that $A$ does not share part of its boundary with that of $\Omega$, that is $\overline{A}\subset\Omega$.
@@ 1.4. Coexistence solutions in higher dimensions
The indicator function of the set $A$, $\mathbb{1}_{A}$ is defined as usual by $\mathbb{1}_{A}(x)=1$ if $x\in A$, else $\mathbb{1}_{A}(x)=0$ if $x\in\overline{\Omega}\setminus A$. In the boundary conditions, $\nu$ stands for the outward unit normal vector on $\Omega$ or $A$ respectively. In this system, the set $A$ represents the predation zone, while $B:=\Omega\setminus\overline{A}$ is the exclusion zone. The previous section dealt with the case $N=1,\Omega=(0,L),A=(0,a)$ (with a common boundary point at 0). We consider here the more general higher dimensional framework. As usual, we note by $B_{r}(a)$ the open ball of radius $r$ centered at a point $a$ and we write $B_{r}=B_{r}(0)$.
@@
Consider the system (4) with $f$ of bistable type satisfying the previous assumptions. Assume that $\gamma<\alpha$.
There exists $R_{0}$ so that for any $A,\Omega$ and such that $A\subset B_{\rho}\subset B_{\rho+R_{0}}\subset\Omega$, there exists a solution $(u,v)$ of system (4) with $v>0$ in $A$ and $u$ is non constant.
@@
As with Theorem 1.3, Theorem 1.4 shows that predators and prey may coexist provided that the prey have a large enough predator-free exclusion zone. We will see that the proof of this Theorem yields the additional result that if the gap between the predator domain and the edge of the prey domain is sufficiently large, the prey population density away from the predation zone can be made arbitrarily close to 1. That is, the prey population far from the predator domain is relatively unaffected by the predators.
1.5. Related literature
@@
The study of diffusive predator-prey models with the type of spatial heterogeneity introduced via an exclusion zone has developed since the foundational 2006 paper of Du and Shi [13], which introduced the notion of a protection zone as a portion of the prey species’ territory from which the predators are excluded. In this work, it is demonstrated that the protection zone fundamentally alters the system dynamics. They use a Holling Type II predation term and logistic growth for the prey. The existence of a critical size for the protection zone, which is expressed through the principal eigenvalue of the Laplacian, determines whether the prey can persist despite predation. When the protection zone exceeds this threshold, the prey persists there, even under high predator growth rates. On the other hand, smaller protection zones may lead to extinction (they consider a generalist predator species which may survive even without the prey species, hence the prey can be hunted to extinction without a resulting extinction of the predators). This framework established the concept of protection zones and emphasized the delicate interplay between spatial geometry, diffusion, and persistence in reaction-diffusion systems.
@@
Subsequent work [12] has extended the concept of protection zones in several directions. In particular, the impact of the strong Allee effect for the prey is explored in [12]. It is shown again that the refuge provided by a protection zone can prevent “over-exploitation” wherein excessive predation drives both species to extinction. Specifically, it is shown that coexistence is possible if the protection zone exceeds a critical size. The analysis reveals bistability and threshold phenomena absent from the monostable formulation of [13]. This line of investigation is further explored in [36] by considering ratio-dependent predation that accounts for predator interference and developing a more detailed bifurcation analysis. They derive explicit analytical criteria for persistence and extinction, prove global existence and positivity of solutions, and demonstrate as in earlier works that spatially non-uniform solutions can bifurcate from semi-trivial (i.e. prey-only) steady states. Together, these studies establish the joint importance of Allee-type bistability and spatial exclusion in promoting coexistence and reveal how the interplay between nonlinear growth and refuge geometry can generate rich spatial dynamics.
@@
Existing studies generally fall into two classes: monostable prey models emphasizing global stability and uniqueness [30, 55], and bistable or Allee-type models focusing on critical protection zone size and bifurcation [12, 36, 14, 53]. In the context of coupled predator-prey systems, most analyses have relied on local bifurcation theory (typically perturbations from semi-trivial equilibria) to establish the existence of coexistence equilibria [12, 36, 57]. Consequently, the resulting positive equilibria are confined to parameter regimes near the onset of bifurcation, and these approaches offer little information about coexistence far from such thresholds. Missing from the literature is a discussion of steady-state solutions far from and independent of bifurcation points. Also absent is an analytical treatment of the case of large protection zones when predators are confined to an asymptotically small region and a numerical demonstration of solution behaviors far from equilibrium and bifurcations associated with these behaviors. Finally, the focus of previous work is consistently on the effect of protection zones on the prey population, with little emphasis on the consequences for predators.

### v2

@@
Abstract
We propose a reaction–diffusion model of predator–prey interaction in which the predators occupy only a subset of the prey’s territory, leaving a predator-free exclusion zone. Ecological examples include marine protected areas where it is illegal to fish, or buffer zones left between the territories of rival predators. The prey are subject to a strong Allee effect, so excessive predation may lead to the extinction of both species. The exclusion zone mitigates this problem by providing the prey with a refuge in which to proliferate without predation. Thus, paradoxically, a smaller predator territory may be able to support a more substantial population than a larger one. Using a topological degree argument, we show in any dimension that, provided the exclusion zone is large enough, the system possesses spatially heterogeneous coexistence equilibria with positive populations of both species. This result is global in the sense that it does not rely on local bifurcations from semi-trivial stationary states. We also show that, as the predator domain becomes asymptotically small, the total predator population does not vanish and in fact approaches an explicit positive limit. In some cases, the total predator population may actually be maximized in this limit of shrinking predation area. Conversely, we show that as the predator domain becomes large, it may exhibit thresholding behavior, passing suddenly from a regime with coexistence solutions to one in which extinction becomes unavoidable, highlighting the need for careful analysis in the management of predator–prey systems.
2020 Mathematics Subject Classification. Primary 35K57; Secondary 35J57, 35B40, 92D25.
Keywords. reaction–diffusion system, predator–prey model, Allee effect, topological degree, exclusion zone, protection area, thin-domain limits, singular limits
1 Introduction
1.1 The role of exclusion zones in predator–prey systems
@@
A somewhat related phenomenon is to be found in the spatial organization of competing (i.e., mutually hostile) groups of predators. For example, groups of predators such as packs of wolves or coyotes often divide available space into sharply defined and defended regions, leading to territory formation [37]. As a consequence of this division, the interfaces between territories constitute buffer zones into which predators avoid venturing and where prey thrive [28]. It was shown [5] that such territoriality can be beneficial for the overall predator population in spite of casualties from the intraspecific hostility. Thus, in practice, the buffer zones between territories act as exclusion zones in which predators are excluded and prey can proliferate without predation.
@@
Of particular interest is the regime in which the consumer’s domain becomes asymptotically small. It may seem that in this case, the consumer population must vanish as its harvest territory becomes small. However, we shall see that their population approaches a positive limit, and in some circumstances, the smaller the consumer-occupied subdomain becomes, the more consumers can be supported. This counterintuitive result shows the subtle interplay between the geometry of consumer and resource territories and the biological mechanisms governing population dynamics.
1.2 The Predator–Prey model with exclusion zones
@@
$\begin{cases}u_{t}-d_{u}\Delta u=f(u)-\beta\mathbb{1}_{A}uv&x\in\Omega\\ v_{t}-d_{v}\Delta v=-\gamma v+\alpha uv&x\in A\\ \partial_{\nu}u(x,t)=0&x\in\partial\Omega\\ \partial_{\nu}v(x,t)=0&x\in\partial A.\end{cases}$ (1)
where $\beta$ is the predation rate, $\gamma$ is the mortality rate of predators, $\alpha$ is a conversion rate, $d_{u}$ and $d_{v}$ are diffusion coefficients of the prey and predator, respectively, $\Omega\subset\mathbb{R}^{N}$ is the prey domain, $A\subset\Omega$ is the predator domain, and $\nu$ represents the outward-pointing normal vector of the boundary of either domain. Moreover, $A\subset\Omega\subset\mathbb{R}^{N}$ are open, $\alpha,\beta,\gamma>0$, and $f:[0,1]\to\mathbb{R}$ is a smooth function with the following properties:
  1. (F1)
    $f$ is continuous.
  2. (F2)
    There are exactly three roots of $f$: 0, 11, and $\theta\in(0,1)$.
  3. (F3)
    For $x\in(\theta,1)$, $f(x)>0$ and for $x\in(0,\theta)$, $f(x)<0$.
  4. (F4)
    $\int_{0}^{1}f(x)\,dx>0$.
We consider only solutions $u$ and $v$ so that $0\leq u(t,x)\leq 1$ for all $x\in\Omega$ and $0\leq v(t,x)$ for all $x\in A$. Note that these definitions stipulate that $\Omega\setminus A$ is the predator-free exclusion zone (equivalently, the MPA) and that both predators and prey may occur in AA; this will later prove notationally convenient. Figure 1 shows an example of $\Omega$ and AA.
Figure 1: A diagram of $A\subset\Omega$ illustrating the assumptions of Theorem 1.2.
As a first result about the model (1), we prove that the predator population can only persist if the predators’ conversion rate $\alpha$ exceeds their mortality rate:
Proposition $1.$
If $\alpha<\gamma$, then all solutions to (1) with $0\leq u\leq 1$ and $v\geq 0$ satisfy $\lim_{t\to\infty}\|v\|_{L^{1}}=0$.
@@
Since the analysis shows that predators inevitably go extinct when $\alpha<\gamma$, this parameter regime is not biologically relevant for sustained predator–prey interactions, and we therefore restrict attention to the case $\gamma\leq\alpha$.
Sections 1.3-4 focus on analytical results for solutions of this model. Much of this analysis concerns equilibrium solutions to (1). Clearly, there are three homogeneous equilibria, namely $v=0$ while $u\in\{0,\theta,1\}$. But we are most interested in inhomogeneous coexistence equilibrium solutions to (1) in which $0<u<1$ and $v>0$. In Section 1.3, we present the main existence results of this paper, showing that such coexistence equilibria exist in both one and higher dimensions. We prove the existence of coexistence equilibria in Section 2. The resulting existence theorems include as a condition that the predator-free domain $\Omega\setminus A$ must be sufficiently large. In Section 3, we demonstrate how coexistence equilibria may cease to exist when this condition does not hold. We perform asymptotic analyses of inhomogeneous equilibria as AA becomes small (so-called “thin-limit” analysis) in one dimension in Section 4.
@@
  • What is the size (or shape) of the predator domain AA that maximizes the predator population?
@@
1.3 Coexistence solutions in higher dimensions
One of the first questions we may ask about the model (1) is about the existence of coexistence equilibria. We consider the predator–prey model, where predation is limited to a predation zone $A\subset\Omega\subset\mathbb{R}^{N}$ (we are most interested in the cases $N=2,3$). We assume AA to be a smooth subset (connected or not) of $\Omega$. We assume here that AA does not share part of its boundary with that of $\Omega$, that is $\overline{A}\subset\Omega$.
@@ 1.3 Coexistence solutions in higher dimensions
The indicator function of the set AA, $\mathbb{1}_{A}$ is defined as usual by $\mathbb{1}_{A}(x)=1$ if $x\in A$, else $\mathbb{1}_{A}(x)=0$ if $x\in\overline{\Omega}\setminus A$. In the boundary conditions, $\nu$ stands for the outward unit normal vector on $\Omega$ or AA respectively. In this system, the set AA represents the predation zone, while $B:=\Omega\setminus\overline{A}$ is the exclusion zone. As usual, we note by $B_{r}(a)$ the open ball of radius $r$ centered at a point $a$ and we write $B_{r}=B_{r}(0)$.
@@
Consider the system (2) with $f$ of bistable type satisfying the previous assumptions. Assume that $\gamma<\alpha$. There exists $R_{0}$ so that for any $A,\Omega$ and such that $A\subset B_{\rho}\subset B_{\rho+R_{0}}\subset\Omega$, there exists a solution $(u,v)$ of system (2) with $v>0$ in AA and $u$ is nonconstant.
@@
Theorem 1.2 shows that predators and prey may coexist provided that the prey have a large enough predator-free exclusion zone. We will see that the proof of this Theorem yields the additional result that if the gap between the predator domain and the edge of the prey domain is sufficiently large, the prey population density away from the predation zone can be made arbitrarily close to 1. That is, the prey population far from the predator domain is relatively unaffected by the predators.
The condition of Theorem 1.2 that AA is a proper subset of $\Omega$ cannot be dropped in general. As we shall see in Section 3, if $A=\Omega$, then no coexistence equilibrium is possible if $\gamma$ is sufficiently small. In this case, the predators are too effective for their own good. They over-hunt the prey, driving them below their own Allee threshold, leading to extinction of the prey followed by the extinction of the predators.
1.4 Related literature
@@
The study of diffusive predator–prey models with the type of spatial heterogeneity introduced via an exclusion zone has developed since the foundational 2006 paper of Du and Shi [13], which introduced the notion of a protection zone as a portion of the prey species’ territory from which the predators are excluded. In this work, it is demonstrated that the protection zone fundamentally alters the system dynamics. They use a Holling Type II predation term and logistic growth for the prey. The existence of a critical size for the protection zone, which is expressed through the principal eigenvalue of the Laplacian, determines whether the prey can persist despite predation. When the protection zone exceeds this threshold, the prey persists there, even under high predator growth rates. On the other hand, smaller protection zones may lead to extinction (they consider a generalist predator species which may survive even without the prey species, hence the prey can be hunted to extinction without a resulting extinction of the predators). This framework established the concept of protection zones and emphasized the delicate interplay between spatial geometry, diffusion, and persistence in reaction-diffusion systems.
@@
Subsequent work [11] has extended the concept of protection zones in several directions. In particular, the impact of the strong Allee effect for the prey is explored in [11]. It is shown again that the refuge provided by a protection zone can prevent “over-exploitation” wherein excessive predation drives both species to extinction. Specifically, it is shown that coexistence is possible if the protection zone exceeds a critical size. The analysis reveals bistability and threshold phenomena absent from the monostable formulation of [13]. This line of investigation is further explored in [35] by considering ratio-dependent predation that accounts for predator interference and developing a more detailed bifurcation analysis. In [35], the authors derive explicit analytical criteria for persistence and extinction, prove global existence and positivity of solutions, and demonstrate as in earlier work that spatially non-uniform solutions can bifurcate from semi-trivial (i.e., prey-only) steady states. Together, these studies establish the joint importance of Allee-type bistability and spatial exclusion in promoting coexistence and reveal how the interplay between nonlinear growth and refuge geometry can generate rich spatial dynamics.
@@
Existing studies generally fall into two classes: monostable prey models emphasizing global stability and uniqueness [29, 54], and bistable or Allee-type models focusing on critical protection zone size and bifurcation [11, 35, 12, 52]. In the context of coupled predator–prey systems, most analyses have relied on local bifurcation theory (typically perturbations from semi-trivial equilibria) to establish the existence of coexistence equilibria [11, 35, 56]. Consequently, the resulting positive equilibria are confined to parameter regimes near the onset of bifurcation, and these approaches offer little information about coexistence far from such thresholds. Missing from the literature is a discussion of steady-state solutions far from and independent of bifurcation points. Also absent is an analytical treatment of the case of large protection zones when predators are confined to an asymptotically small region and a numerical demonstration of solution behaviors far from equilibrium and bifurcations associated with these behaviors. Finally, the focus of previous work is consistently on the effect of protection zones on the prey population, with little emphasis on the consequences for predators.